Fermi National Laboratory


Accelerator Update

Accelerator Update

The integrated luminosity for last week was 4.1 pb-1.

Wednesday September 3
The day shift began with the Antiproton Source (Pbar) stacking and with Tevatron experts conducting studies.

TeV experts completed their studies at 10 AM and Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 10:31 AM.

At 12:00 PM, the TeV suffered a sector A1 quench just after the dampers were turned on. Pbar resumed stacking. TeV experts began a study period to determine what happened.

Operators made the D-zero (D0) and the Collider Detector Facility (CDF) ready for access at 12:25 PM. D0 came out of access at 1:22 PM and CDF came out at 3 PM.

Operations loaded a 36x0 TeV study-store at 4:11 PM.

A Linac RF station (LRF2) tripped off at 4:36 PM. Operators reset the station and had it back on ten minutes later.

TeV expert completed their damper studies at 7 PM. They decided that the dampers should not be used for the next store.

Operations put the TeV into shot setup at 7:30 PM. At 9:41 PM, Operations established store 2985 with an initial luminosity of 3.4E31.

At 11:38 PM, the TEL output went to zero. Operators investigated and then called in an expert.

Thursday September 4
The midnight shift began with Operations monitoring store 2985, with Pbar stacking, and with Main Injector (MI) experts conducting 2.5MHz studies.

The TEL experts reported that he replaced a power supply at 1:30 AM.

The MI experts completed their 2.5MHz studies at 2:24 AM, but then started a slipstacking study. This study ended at 7:32 AM.

The Cryo Coordinator informed Operations at 9:50 AM that the TeV sector A3 wet engine was suffering serious problems. It will have to be worked on as soon as the store ends.

A Pbar expert began a multiwire study on the P1 and P2 transfer lines at 10:24 AM.

Operations halted stacking at 3:02 PM to prepare MI for an access.

Operations terminated store 2985 at 3:04 PM.

Cryo technicians began work on the A3 wet engine at 3:10 PM. They had the job finished by 4:28 PM.

The MI access was over by 4:31 PM.

Operations put the TeV through a dry squeeze at 4:49 PM.

Pbar resumed stacking at 5:21 PM.

Operations put the TeV into shot setup at 6:45 PM. At 9:15 PM, Operations established store 2988 with an initial luminosity of 3.6E31. Pbar resumed stacking.

At 9:25 PM, Operations reported that the TeV luminosities were dropping rapidly. The problem began when operators turned on the TEL. An expert was called in to investigate.

The TEL expert corrected a triggering problem at 10:31 PM.

Friday September 5
The midnight shift began with Operations monitoring store 2988 and with Pbar experts conducting studies.

At 4:42 AM, Pbar experts completed their studies. Unfortunately, stacking did not resume due to a problem with D:ISEPV. After contacting an expert, operators cycled a breaker. This helped a little, but stacking was very poor. Operators cycled several other supplies, and this also helped a little. By 5:40 AM, experts were tuning. At 6:10 AM, Debuncher RF cavities (DRF1) were warming up to normal operating temperatures.

MI experts began a study period at 5:37 AM.

At 6:13 AM, the TeV RF went unstable for a moment due to TRF1. Operators turned this station off.

A Linac RF station (LRF5) tripped off at 6:24 AM. Operators had the station back on in five minutes.

Today's Plan
The plans for today are to allow an end-of-store study period to begin around 10 AM and then terminate the store around noon. Operations will put the TeV into shot setup. The plan for the weekend is to establish two stores between now and Sunday. Operations will terminate the last store on Sunday at 4 PM. The seven to ten-week shutdown begins at 8 AM on Monday morning, September 8, 2003.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/5/2003   email Fermilab